Petrol​ Ibérico, a European gas​ company, is borrowing ​$500,000,000 via a syndicated eurocredit for six years at 70 basis points over LIBOR. LIBOR for the loan will be reset every six months. The funds will be provided by a syndicate of eight leading investment​ bankers, which will charge​ up-front fees totaling 1.3​% of the principal amount. What is the effective interest cost for the first year if the annual LIBOR is 4.10​% during the first six months and 4.50​% during the second six months.
Answer #1

1st 6 months LIBOR is 4.1%; Borrowing cost is 4.1%+.7%=4.8%

2nd 6 months LIBOR is 4.5%; Borrowing cost is 4.5%+.7%=5.2%

If 100$ were the loan amount,

Amount at end of 1st 6 months is 100(1+4.8%)^0.5=102.37

Amount at end of 2nd 6 months is 102.37(1+5.2%)^0.5=104.99

Effective rate is (104.99-100)/100=4.99%

In addition 1.3% is up front fee

Hence total interest cost is 4.99%+1.3%=6.29%
Answer #2

Its wrong bro

source: Trust me bro
answered by: anonymous